OpenStudy (anonymous):

Which of the following statements represents the lowest amount? Salary before an 18% pay raise to $45,000 Income on a $45,000 salary after 18% income tax Salary before an 18% pay cut to $45,000 $45,000 plus 18% commission on $45,000

OpenStudy (anonymous):

???

OpenStudy (anonymous):

salary before an 18% raise=(100-18)/100*45000= 82/100*45000=$36900 B)income tax=18/100*45000=8100

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20110125183002AANg0Am medal plz :)

OpenStudy (anonymous):

thxx

OpenStudy (anonymous):

anytime !